OpenStudy (anonymous):

Dad is 4 times as old as his son Jim. In 10 years, Dad's age will be 20 years more than twice Jim's age. How old is Jim? If D = dad's age now and J = Jim's age now, which of the following systems could be used to solve the problem? D = 4J and D + 10 = 2J + 20 D = 4J and D + 10 = 2(4J + 10) + 20 D = 4J and D + 10 = 2(J + 10) + 20 A?

OpenStudy (anonymous):

Right now, let Jim's age = x So, Dad's age is 4x In ten years from now, Jim's age: x + 10 Dad's age is 4x + 10 So according to the problem, in 10 years, 4x + 10 = 2(x+10) + 20

OpenStudy (anonymous):

I would use B: D = 4J and D + 10 = 2(4J + 10) + 20